# Resource Allocation for Twin Maintenance and Computing Task Processing in Digital Twin Vehicular Edge Computing Network

### v1

@@ V-C Performance Evaluation
In Figure 4, we compared the average rewards of MARL-CSTC algorithm with SAC algorithm, PPO algorithm, and random algorithm. As the number of iterations increases, the average rewards of the other three algorithms, except for the random algorithm, will eventually converge to a larger value. Specifically, when the number of iterations is small, due to lack of experience, the total amount of resources requested by the agent from the server may exceed the computational resources of the server itself. In addition, agents may also be punished for violating constraints (16b) and (16c) due to improper resource allocation strategies in the early stage. As the number of iterations gradually increases, the agent is able to explore appropriate resource allocation strategies within the total computational range of the VEC server, resulting in a gradual convergence of reward values. The fluctuation after convergence is due to the dynamic randomness of the environment, which can affect learning. Among the three algorithms, the average converged reward value by the MARL algorithm is greater than that of the SAC algorithm and PPO algorithm. This is because compared to algorithms such as SAC and PPO, our proposed MARL-CSTC algorithm can allocate server computing resources reasonably for twin maintenance and computing task processing based on current number of vehicles, channel resources, transmission power, and other information. The reason why the random algorithm cannot converge is because it first distributes the total resources evenly and then randomly allocates the number of resources, so its average reward value always fluctuates around a certain value, but it is still lower than the average reward value of the final convergence of MARL-CSTC.
Figure 5 shows the resource utilization rate and the ratio of resource conversion to utility under different algorithms and vehicle numbers. From Figure 5 (a), it can be seen that MARL-CSTC achieves the highest resource utilization rate. This is because MARL-CSTC can fully schedule the computing resources, i.e. CPU frequency, of the VEC server based on the twin maintenance and vehicular task requirements of all vehicles. On the contrary, PPO and SAC are based on local information of a single vehicle for resource scheduling and sharing with other vehicles, thus unable to fully utilize server computing resources. Due to the random nature of twin maintenance and vehicular task requirements for individual vehicles, as well as changes in the number of vehicles, the previous resource scheduling plan is often not applicable and needs to be redesigned. Therefore, the resource utilization rates of PPO and SAC will fluctuate. From Figure 5 (b), it can be seen that as the number of vehicles increases, the ratio of computing resources to utility by each algorithm will decrease, while MARL-CSTC can still achieve higher conversion ratios than PPO and SAC. As mentioned earlier, due to its ability to rely on global information, MARL-CSTC is less affected by changes in the number of vehicles. In addition, MARL-CSTC will prioritize meeting the twin maintenance and vehicular task requirements of each vehicle in its specific allocation plan, thus achieving a higher conversion rate.
Figure 6 shows the maximum utility that a single vehicle can achieve under different influencing factors. From Figure 6 (a), it can be seen that when the total number of vehicles is small, a single vehicle can achieve higher resource utility. As the number of vehicles increases, the resource utility of each vehicle will decrease accordingly.This is because an increase in the number of vehicles will require the server to maintain more digital twin models and handle more computing tasks. Therefore, the computing resources allocated to individual vehicles will be reduced. However, the performance achieved under the MARL-CSTC algorithm is still higher than other algorithms. This is because according to Figure 5(a), MARL-CSTC has higher resource utilization than SAC, PPO, and Random, which means that MARL-CSTC has more assignable computing resources. From Figure 6(b), when the total number of vehicles is 5, as the transmission power of the vehicles increases, the utility achieved by each algorithm will also increase. An increase in transmission power means that under the same channel resource conditions, the transmission rate will also increase, which in turn will reduce the possibility of exceeding the time limit for twin maintenance and computing tasks. Therefore, computing resources originally used to meet the time limit conditions can be allocated to achieve higher utility.
(a) Twin maintenance delay
(b) Vehicular computing task processing delay
Figure 7: Comparison of two types of delays under different algorithms:$p_{i}^{max}$=200mW, $N$=5
Figure 7 examines the latency generated by each algorithm while performing twin maintenance and computing task processing simultaneously. We can observe that as the number of vehicles in the environment increases, the latency generated by each algorithm also increases. Specifically, when the number of vehicles is 5, except for the random algorithm, the other three algorithms can achieve lower latency while satisfying time constraints, and the MARL algorithm achieves the minimum latency. As the number of vehicles gradually increases, except for the MARL-CSTC algorithm, which can reduce latency under time constraints in twin maintenance and computational task processing, the other three types of algorithms cannot simultaneously handle these two tasks. According to Figure 5, this is because on the one hand, MARL-CSTC is used for more allocatable computing resources, and on the other hand, it has a higher resource conversion ratio, which prioritizes meeting the time constraints of twin maintenance and onboard tasks, and on this basis, the remaining computing resources are re allocated. In addition, SAC and PPO algorithms tend to favor twin maintenance as the number of vehicles gradually increases, while random algorithms tend to favor computing tasks.
@@ VI Conclusions
This paper considered the server resource allocation problem of twin maintenance and computing task processing simultaneously under edge computing for a digital twin scenario with single server and multiple vehicles, and constructed it as an optimization problem to maximize the resource utility of each vehicle. To address the non convexity of the problem, we first adopted multi-agent MDP to rephrase the problem and proposed the MADRL-CSTC algorithm to obtain the optimal resource allocation. Experimental results have shown that compared with other algorithms, MADRL-CSTC can maximize the resource utility of each vehicle under constraint conditions. The conclusion is summarized as follows:
@@
  • Our method provides an effective resource allocation for each vehicle within the total server resource range, under the condition of twin maintenance and computing task deadlines.

### v2

@@ V-C Performance Evaluation
Figure 4 compares the average rewards of the different algorithms. As the number of iterations gradually increases, the other three algorithms, except for the random algorithm, are able to increase from small to large values. This shows the effectiveness of DRL. In addition, TMTPRCS achieves higher reward values than SAC, PPO and Random algorithms. Compared to SAC, PPO, Random and MATD3 algorithms, the reward value achieved by TMTPRCS after convergence is $6.42\%$, $11.15\%$, $37.56\%$ and $83.42\%$ higher than them respectively. In a fully competitive setting, our method works well to obtain the optimal policy whereas MATD3 converges to a locally optimal solution due to insufficient exploration.
Figure 5 compares the resource utilization of different algorithms for different number of vehicles. As can be seen in Fig. 5.a, TMTPRCS always maintains a high resource utilization compared to SAC ,PPO, and MATD3. On the contrary, SAC, PPO and MATD3 only gradually increase its resource utilization when the number of vehicles gradually increases. For TMTPRCS, the reason that the resource utilization is higher than the other numbers when the number of vehicles is 6 is that the algorithm achieves the optimal resource allocation balance in this case, i.e., there are neither idle nor overloaded resources. Specifically, when the number of vehicles is lower than 6, the number of twin models to be maintained in the server is lower, which results in less resource demand and, therefore, slightly lower utilization of resources. When the number of vehicles is higher than 6, the demand for resources increases, but the total resources in the server are limited, thus leading to a decrease in the overall utilization of resources. This numerical result also indicates that when the resources of twin model are limited, there exists an optimal served number of vehicles, which needs to be determined in theory in the future. In fact, TMTPRCS is able to collect global information about all vehicles and the requirements corresponding to twin maintenance tasks and computation tasks due to the centralized training mentioned in Fig.4. On the contrary, PPO and SAC only make decisions based on the local information of a single vehicle and share it directly to other vehicles. In addition, MATD3, as described in Fig. 4, it converges to a locally optimal solution due to under-exploration, thus leading to under-utilization of resources under this strategy. In order to visualize the resources that can be dispatched by the vehicles when performing twin maintenance and task processing, we present Fig. 5.b and Fig. 5.c. It can be seen that the amount of resources that can be scheduled by TMTPRCS is higher than that of the other three.
Figure 6 illustrates the maximum efficiency that can be achieved for each vehicle at different numbers of vehicles. As the number of vehicles gradually increases, the effectiveness achieved per vehicle decreases. Obviously, an increase in the number of vehicles implies an increase in the demand for twin maintenance and task processing. Due to the resource competition between vehicles, vehicles will always request more resources to avoid the failure of these two tasks. However, the total resources of the server remain the same, so under different algorithm scheduling, the resources that each vehicle can request will also decrease, which in turn decreases the performance. In addition, TMTPRCS can better cope with competitive situations with multiple vehicles because it can be based on global information.
(a) Delay of twin maintenance
(b) Delay of vehicular computing task processing
Figure 7: Comparison of two types of delays under different algorithms:$p_{i}$=200mW
Figure 7 compares the twin maintenance delay and task processing delay of different algorithms for different number of vehicles. It can be seen that both types of delays gradually increase with the number of vehicles. Specifically, in Fig. 7.a, all four types of algorithms, except for the random algorithm, meet the deadline for the twin maintenance task. Here, the reason for the slower delay growth when the number of vehicles increases from 5 to 6 is that the algorithm is able to allocate resources more efficiently in this case, reducing resource competition. Specifically, the delay growth rate in this range is only $8.91\%$, significantly lower than the previous stage. In contrast, when the number of vehicles increases from 4 to 5, the demand for resources rises while the allocation is not yet optimal, resulting in a high delay growth rate of $104.41\%$, although the total delay remains lower than that observed at 5 and 6 vehicles. Similarly, as the number of vehicles increases from 6 to 7, the resource demand continues to rise, and competition intensifies, leading to a significant $78.16\%$ increase in delay. This trend highlights the limits of the algorithm’s resource management under heavier loads. Nevertheless, despite these variations in delay growth, the overall latency remains within acceptable bounds. And in Fig. 7.b, all five classes of algorithms satisfy the maximum time tolerance for the computation task. In addition, TMTPRCS always achieves lower latency compared to the other three classes of algorithms. This is due to the fact that on the one hand, according to Fig. 5, TMTPRCS has a higher resource utilization and thus more schedulable resources, and on the other hand, according to what is described in Fig. 6, TMTPRCS is able to better deal with resource competition under multiple vehicles. From the two figures, on the one hand, it can be seen that the MATD3 algorithm has the highest task processing latency due to the low resource utilization of the MATD3 algorithm and on the other hand, the locally optimal solution achieved by MATD3 has made it invest its resources in reducing the twin maintenance delay.
@@ VI Conclusions
This paper considered the server resource allocation problem of twin maintenance and computing task processing simultaneously under edge computing for a digital twin scenario with single server and multiple vehicles, and constructed it as an optimization problem to maximize the resource utility of each vehicle. In the future, we will conduct a more in-depth study. In multi-server scenarios, servers can share resource and load data to collaborate. Dynamic load balancing can help assign tasks in real time. In this setting, When TMTPRCS shifts from centralized to decentralized training, it brings partial observability challenges to the critic network. These can be mitigated through joint updates, global state aggregation, or hierarchical coordination. In practical deployments, as vehicles move across base stations, their digital twin models need real-time migration. We aim to enhance TMTPRCS with trajectory-aware resource pre-allocation, lightweight state transfer, and decentralized edge server coordination to support mobility. In addition, while TMTPRCS currently assumes offline centralized training, future work will focus on adapting the algorithm for real-world edge deployments. This includes exploring model compression techniques, lightweight network architectures, and shared policy strategies to reduce computational and memory overhead on edge servers. At the same time, comparisons with classical reinforcement learning methods are insufficient, so we will introduce comparisons with more advanced methods in future work.
The conclusion is summarized as follows:
@@
  • Our approach enables the development of rational resource request policies for each vehicle under twin maintenance and vehicular tasks deadlines and total resource constraints.